# Identification and estimation of mediational effects of longitudinal modified treatment policies

### v1

@@ Abstract
We demonstrate a comprehensive semiparametric approach to causal mediation analysis, addressing the complexities inherent in settings with longitudinal and continuous treatments, confounders, and mediators. Our methodology utilizes a nonparametric structural equation model and a cross-fitted sequential regression technique based on doubly robust pseudo-outcomes, yielding an efficient, asymptotically normal estimator without relying on restrictive parametric modeling assumptions. We are motivated by a recent scientific controversy regarding the effects of invasive mechanical ventilation (IMV) on the survival of COVID-19 patients, considering acute kidney injury (AKI) as a mediating factor. We highlight the possibility of “inconsistent mediation,” in which the direct and indirect effects of the exposure operate in opposite directions. We discuss the significance of mediation analysis for scientific understanding and its potential utility in treatment decisions.
Keywords— mediation, machine learning, longitudinal, treatment policies
@@ 1 Introduction
Even so, the formulation of an estimand of interest in a mediation analysis is not always a simple task. The so-called “natural” mediation effects, where one considers an intervention that partially “blocks” the effect of an exposure by assigning mediator values to their counterfactual values under a fixed exposure level in order to isolate “direct” and “indirect” effects, have been well-studied, and conditions for their identification have been given (Robins and Greenland, 1992; Pearl, 2001). However, those identification conditions are quite strong and unlikely to be satisfied in many instances of real data-generating processes (Avin et al., 2005; Miles et al., 2015). To analyze mediation under more reasonable assumptions, alternative estimands based on stochastic draws from the mediator distribution have been proposed and labeled “interventional effects” (VanderWeele et al., 2014).
Mediation analysis in longitudinal settings is particularly challenging; some existing approaches have relied on parametric model specification (VanderWeele and Tchetgen, 2017; Tai et al., 2022), which leads to a lack of robustness. Zheng and van der Laan (2017) provided a robust machine-learning approach for the estimation of mediation effects with time-varying treatments, confounders, and mediators. This line of research was advanced by Díaz et al. (2023b) which allowed the (interventional) direct effect to include pathways through intermediate confounders. This formulation is desirable if the goal of mediation analysis is to separate effects according to whether they operate through the mediator or not, which is the point of view of the present article.
@@
We synthesize these various strands of the statistical literature into a new strategy for mediation analysis in longitudinal settings. Specifically, we consider the direct and indirect effects of longitudinal modified treatment policies under interventional mediator distributions, which have not been considered in previous literature. The methods we propose are motivated by an issue that was relevant in the early landscape of COVID-19 management: the use of invasive mechanical ventilation (IMV) in cases of acute respiratory distress syndrome (ARDS). ARDS is a form of acute hypoxic respiratory failure, which is one of the key time-varying features of a patient’s course with severe COVID (see, e.g., Wang et al. (2021)). In the beginning stages of the pandemic, IMV was employed early, or at a lower supplemental oxygen delivery threshold, to manage ARDS, but its usage has been associated with several iatrogenic risks such as ventilator-associated pneumonia (Wicky et al., 2021) and barotrauma (Shrestha et al., 2022). Of interest to the present study is acute kidney injury (AKI), a critical condition that complicates ICU stays and is associated with mortality (Vemuri et al., 2022). As more data emerged and practitioners became more familiar with the pace of disease progression, guidelines shifted towards delaying intubation due to high mortality rates among mechanically ventilated patients and potential secondary complications. To navigate these treatment dynamics, we aim to elucidate the direct and indirect effects of IMV on patient mortality with respect to the AKI pathway. This application offers a compelling case for the use of our methods in understanding the effects of time-varying treatments and the causal pathways involved. We emphasize the phenomenon of “inconsistent mediation” (a situation in which the direct and indirect effects are qualitatively in opposite directions). In the current setting, this would involve (on average) IMV causing death through AKI while (on average) preventing death through other mechanisms. This dichotomy may be relevant for research and clinical practice, if interventions could be in place to mitigate IMV-induced AKI. While this is an active area of research, potential interventions include the administration of diuretics to modulate intravascular volume status (Grams et al., 2011; Glassford and Bellomo, 2011) and pharmacological options as explored by Pickkers et al. (2022). Finally, we also investigate the relationship between baseline covariates and the magnitude of these direct and indirect effects; such relationships can suggest biological pathways as well as potentially inform future treatments.
@@ 2.1 Notation
Our notation largely mirrors that of Díaz et al. (2021) and Díaz et al. (2023b). We consider data $X_{1},\ldots,X_{n}$, an i.i.d. sample from a distribution $\mathsf{P}$. The data are longitudinal, with $X=\\ (L_{1},A_{1},Z_{1},M_{1},L_{2},\ldots,A_{\tau},Z_{\tau},\allowbreak M_{\tau},L_{\tau+1})$, where the variables $A$ are the exposure of interest, $L$ are covariates, $M$ are mediators, and $Z$ are intermediate confounders. An intermediate confounder is a variable affected by exposure and which affects the mediator and outcome; see VanderWeele et al. (2014). We let $Y=L_{\tau+1}$ denote the outcome of interest. As in our illustrative application, we will assume that the mediator takes values in a finite set; the supports of all other variables are unrestricted.
@@ 2.2 Structural equation model and definition of causal effects
We consider causal effects defined within a nonparametric structural equation model, which consists of exogenous and endogenous variables and unknown deterministic functions governing the causal relations between them, together with independence assumptions on the exogenous variables (Pearl, 2000). We allow for arbitrary causal relations among observed variables that respect their time-ordering; this can be visualized in the directed acyclic graph (DAG) given in Figure 1, reproduced from Díaz et al. (2023b).
@@
To formalize the dependence structure, we posit functions $f$ such that for each $t$, we have $A_{t}=f_{A,t}(H_{A,t},U_{A,t})$, $Z_{t}=f_{Z,t}(H_{Z,t},U_{Z,t})$, $M_{t}=f_{M,t}(H_{M,t},U_{M,t})$, and $L_{t}=f_{L,t}(H_{L,t},U_{L,t})$. These unobserved “exogenous” variables $U=(U_{A,t},U_{Z,t},U_{M,t},U_{L,t},U_{Y}:t\in\{1,\ldots,\tau\})$. In the general model, there are no restrictions on the distribution of $U$, but in Section 3, they will be subject to assumptions to enable identification of the estimands of interest.
@@
For interventions on the treatment variable, we will consider longitudinal modified treatment policies (LMTPs) (Díaz et al., 2021), which are hypothetical interventions defined in terms of a sequence of functions $\mathbbm{d}=(\mathbbm{d}_{1},\ldots,\mathbbm{d}_{\tau})$, where $\mathbbm{d}_{t}$ is a function $\mathbbm{d}_{t}(a_{t},h_{A,t},\epsilon_{t})$ that depends on a treatment value $a_{t}$, a history value $h_{A,t}$, and possibly a randomizer $\epsilon_{t}$. The intervention is defined by removing the equation $A_{t}=f_{A_{t}}(H_{t},U_{A,t})$ sequentially (from $t=1$ to $t=\tau$) from the structural model, and replacing it with evaluations of the functions $\mathbbm{d}_{t}$ as follows. At the first timepoint, the LMTP assigns the exposure as a new random variable $A_{1}^{\mathbbm{d}}=\mathbbm{d}_{t}(A_{1},H_{A,1},\epsilon_{1})$. This generates a counterfactual treatment $A_{2}(\bar{A}_{1}^{\mathbbm{d}})=f_{A_{2}}(H_{A,2}(\bar{A}_{1}^{\mathbbm{d}}),U_{A,1})$ that would have been observed if the intervention had been discontinued right before $t=2$. Then, the intervention at $t=2$ assigns treatment as $A_{2}^{\mathbbm{d}}=\mathbbm{d}_{2}(A_{2}(\bar{A}_{1}^{\mathbbm{d}}),H_{A,2}(\bar{A}_{1}^{\mathbbm{d}}),\epsilon_{2})$. In general, an intervention that sets the exposure values up to time $t-1$ to $\bar{A}_{t-1}^{\mathbbm{d}}$ generates counterfactual variables $L_{t}(\bar{A}_{t-1}^{\mathbbm{d}})=f_{L_{t}}(A_{t-1}^{\mathbbm{d}},H_{t-1}(\bar{A}_{t-2}^{\mathbbm{d}}),U_{L,t})$, $M_{t}(\bar{A}_{t-1}^{\mathbbm{d}})$, and $Z_{t}(\bar{A}_{t-1}^{\mathbbm{d}})$. In addition, $A_{t}(\bar{A}_{t-1}^{\mathbbm{d}})=f_{A_{t}}(H_{t}(\bar{A}_{t-1}^{\mathbbm{d}}),U_{A,t})$, $A_{t}(\bar{A}_{t-1}^{\mathbbm{d}})$, termed the “natural value of treatment” (Richardson and Robins, 2013; Young et al., 2014), “represents the value of treatment that would have been observed at time $t$ under an intervention carried out up until time $t-1$ but discontinued thereafter” (Díaz et al., 2021). The intervention at time $t$ is given by $A_{t}^{\mathbbm{d}}=\mathbbm{d}_{t}(A_{t}(\bar{A}_{t-1}^{\mathbbm{d}}),H_{A,t}(\bar{A}_{t-1}^{\mathbbm{d}}),\epsilon_{t})$.
@@
The intervention on $\bar{M}$ is defined in terms of a stochastic intervention. Specifically, let $M_{t}(\bar{A}^{\mathbbm{d}})=f_{M,t}(A_{t}^{\mathbbm{d}},H_{M,t}(\bar{A}_{t}^{\mathbbm{d}}),U_{M,t})$ denote the mediator at time $t$ observed under the above LMTP, and let $\bar{J}(\bar{A}^{\mathbbm{d}})$ denote a random draw from the distribution of $\bar{M}(\bar{A}^{\mathbbm{d}})$, potentially conditioning on baseline variables $L_{1}$. We consider interventions where $\bar{M}$ is replaced by this random draw.
@@
A causal “effect” is typically conceived as some difference in outcome distributions under different interventions. To simplify notation throughout, we will consider two user-specified, fixed interventions $\mathbbm{d}^{\star}=(\mathbbm{d}^{\star}_{1},\ldots,\mathbbm{d}^{\star}_{\tau})$ and $\mathbbm{d}^{\prime}=(\mathbbm{d}^{\prime}_{1},\ldots,\mathbbm{d}^{\prime}_{\tau})$, though the generalization of the theory to arbitrary $\mathbbm{d}$ is immediate. We define the randomized modified treatment policy effect as $\mathsf{E}[Y(\bar{A}^{\mathbbm{d}^{\prime}},\bar{J}(\bar{A}^{\mathbbm{d}^{\prime}}))-Y(\bar{A}^{\mathbbm{d}^{\star}},\bar{J}(\bar{A}^{\mathbbm{d}^{\star}}))]$ and decompose it into an interventional direct effect
$DE=\mathsf{E}[Y(\bar{A}^{\mathbbm{d}^{\prime}},\bar{J}(\bar{A}^{\mathbbm{d}^{\star}}))-Y(\bar{A}^{\mathbbm{d}^{\star}},\bar{J}(\bar{A}^{\mathbbm{d}^{\star}}))]$ (1)
@@
$IE=\mathsf{E}[Y(\bar{A}^{\mathbbm{d}^{\prime}},\bar{J}(\bar{A}^{\mathbbm{d}^{\prime}}))-Y(\bar{A}^{\mathbbm{d}^{\prime}},\bar{J}(\bar{A}^{\mathbbm{d}^{\star}}))],$ (2)
which sum to the “overall” or “total” effect comparing $\mathbbm{d}^{\prime}$ to $\mathbbm{d}^{\star}$. Thus, the task at hand is to estimate expressions of the form $\mathsf{E}[Y(\bar{A}^{\mathbbm{d}^{\prime}},\bar{J}(\bar{A}^{\mathbbm{d}^{\star}}))]$, for general $\mathbbm{d}^{\star}$ and $\mathbbm{d}^{\prime}$.
@@ 3.1 Assumptions and sequential regression
In what follows, it will be necessary to reference variable histories under interventions. For example, we let $H_{L,t}^{\prime}=(\bar{L}_{t-1},\bar{M}_{t-1},\bar{Z}_{t-1},\bar{A}_{t-1}^{d^{\prime}})$. Intervened histories for other variables are defined analogously, as are counterfactual histories under the other intervention $\mathbbm{d}^{\star}$, denoted with $H^{\star}$. Similarly, to ease notation, we let $A_{t}^{\prime}=\mathbbm{d}_{t}^{\prime}(A_{t},H_{A,t},\epsilon_{t})$, with $A_{t}^{\star}$ defined similarly.
To identify the parameter, $\mathsf{E}[Y(\bar{A}^{\mathbbm{d}^{\prime}},\bar{J}(\bar{A}^{\mathbbm{d}^{\star}}))]$, the following assumptions, which combine elements of the identification assumptions from Díaz et al. (2021) and Díaz et al. (2023b), are sufficient. Note that A1 involves future histories rather than single timepoints only.
A1Conditional exchangeability of treatment and mediator.
@@
    $U_{A,t}\mbox{$\perp\!\!\!\perp$}(\underline{U}_{Z,t},\underline{U}_{M,t},\underline{U}_{L,{t+1}},\underline{U}_{A,t+1})\mid H_{t}$ for all $t\in\{1,\ldots,\tau\}$;
@@
    $U_{M,t}\mbox{$\perp\!\!\!\perp$}(\underline{U}_{L,{t+1}},\underline{U}_{A,t+1}),\underline{U}_{Z,{t+1}},\mid H_{t}$ for all $t\in\{1,\ldots,\tau\}$.
A2Positivity of treatment and mediator assignment.
@@
    If $(a_{t},h_{A,t})\in\mathop{\mathrm{supp}}\{A_{t},H_{A,t}\}$ then $(\mathbbm{d}(a_{t},h_{A,t}),h_{A,t})\in\mathop{\mathrm{supp}}\{A_{t},H_{A,t}\}$ for $t\in\{1,\ldots,\tau\}$;
@@
A3Piecewise-smooth invertibililty of treatment modification.
If $A_{t}$ is not a discrete random variable, the conditional support of $A_{t}$ given $H_{A,t}=h_{A,t}$ admits a partition into subintervals where the restrictions of $\mathbbm{d}^{\prime}(\cdot,h_{a,t})$ and $\mathbbm{d}^{\star}(\cdot,h_{a,t})$ to each subinterval have differentiable inverse functions.
Our identification strategy makes use of a representation of longitudinal causal estimands by means of “sequential regression” as proposed by Bang and Robins (2005) and further investigated by van der Laan and Gruber (2012); Luedtke et al. (2017), and Rotnitzky et al. (2017). This approach considers a recursive regression of outcomes (or pseudo-outcomes) onto previous variables, starting at $t=\tau$ and continuing backward to $t=0$. As in Díaz et al. (2023b), we split the recursion over the two non-interventional variables $L$ and $Z$ as follows. We consider fixed, pre-specified interventions $\mathbbm{d}^{\prime}$, $\mathbbm{d}^{\star}$, and $\bar{m}$. Starting with, $\mathsf{Q}_{Z,\tau+1}=Y$, define
@@
When the meaning is clear, noting that the only non-random component of the functions $\mathsf{Q}$ are the user-specified $\mathbbm{d}^{\prime},\mathbbm{d}^{\star}$, and $\bar{m}$, we may use the abbreviations $\mathsf{Q}_{L,t}^{\prime}(\underaccent{\bar}{m}_{t})=\mathsf{Q}_{L,t}(\bar{H}_{M,t},\underaccent{\bar}{m}_{t})$, $\mathsf{Q}_{Z,t}^{\prime}(\underaccent{\bar}{m}_{t})=\mathsf{Q}_{Z,t}(A^{{}^{\prime}}_{t},\bar{H}_{A,t},\underaccent{\bar}{m}_{t})$, and $\mathsf{Q}_{M,t}^{\star}(\underaccent{\bar}{m}_{t})=\mathsf{Q}_{M,t}(A_{t}^{\star},\bar{H}_{A,t},\underaccent{\bar}{m}_{t})$. We give the following identification result, which is a generalization of Theorem 1 of Díaz et al. (2023b).
@@
Under Assumptions A1 and A2, the interventional parameter $\theta=\mathsf{E}[Y(\bar{A}^{\mathbbm{d}^{\prime}},\bar{J}(\bar{A}^{\mathbbm{d}^{\star}}))]$ is identified as $\theta=\sum_{\bar{m}}\varphi(\bar{m})\lambda(\bar{m})$, where $\varphi(\bar{m})=\mathsf{Q}_{L,0}^{\prime}(\bar{m})$ and $\lambda(\bar{m})=\mathsf{Q}_{M,0}^{\star}(\bar{m})$.
Proof See Supplement Section S2. ∎

### v3

@@ Abstract
We demonstrate a comprehensive semiparametric approach to causal mediation analysis, addressing the complexities inherent in settings with longitudinal and continuous treatments, confounders, and mediators. Our methodology utilizes a nonparametric structural equation model and a cross-fitted sequential regression technique based on doubly robust pseudo-outcomes, yielding an efficient, asymptotically normal estimator without relying on restrictive parametric modeling assumptions. We are motivated by a recent scientific controversy regarding the effects of invasive mechanical ventilation (IMV) on the survival of COVID-19 patients, considering acute kidney injury (AKI) as a mediating factor. We highlight the possibility of “inconsistent mediation,” in which the direct and indirect effects of the exposure operate in opposite directions. We discuss the significance of mediation analysis for scientific understanding and its potential utility in treatment decisions.mediation, machine learning, longitudinal, treatment policies
@@ 1 Introduction
Even so, the formulation of an estimand of interest in a mediation analysis is not always a simple task. The so-called “natural” mediation effects, where one considers an intervention that partially “blocks” the effect of an exposure by assigning mediator values to their counterfactual values under a fixed exposure level in order to isolate “direct” and “indirect” effects, have been well-studied, and conditions for their identification have been given (Robins and Greenland, 1992; Pearl, 2001). However, those identification conditions are quite strong and unlikely to be satisfied in many instances of real data-generating processes (Avin et al., 2005; Miles et al., 2015). In particular, they require the assumption of “cross-world counterfactual independence,” where counterfactual outcomes downstream of mutually exclusive counterfactual events are independent, which cannot be tested even experimentally. To analyze mediation under more reasonable assumptions, alternative estimands based on stochastic draws from the mediator distribution have been proposed and labeled “interventional effects” (VanderWeele et al., 2014).
Mediation analysis in longitudinal settings is particularly challenging; some existing approaches have relied on parametric model specification (VanderWeele and Tchetgen, 2017; Tai et al., 2022), which leads to a lack of robustness. Zheng and van der Laan (2017) provided a robust machine-learning approach for the estimation of mediation effects with time-varying treatments, confounders, and mediators. This line of research was advanced, for example, by Díaz et al. (2023b) which allowed the (interventional) direct effect to include pathways through intermediate confounders, which allows the indirect effect to isolate precisely those operating through the mediator and not other variables.
@@
We synthesize these various strands of the statistical literature into a new strategy for mediation analysis in longitudinal settings, following Díaz and Hejazi (2020) and Hejazi et al. (2023), who studied mediation analysis for single-timepoint stochastic interventions. Specifically, we consider the direct and indirect effects of longitudinal modified treatment policies under interventional mediator distributions, which have not been considered in previous literature. The methods we propose are motivated by an issue that was relevant in the early landscape of COVID-19 management: the use of invasive mechanical ventilation (IMV) in cases of acute respiratory distress syndrome (ARDS). ARDS is a form of acute hypoxic respiratory failure, which is one of the key time-varying features of a patient’s course with severe COVID (see, e.g., Wang et al. (2021)). In the beginning stages of the pandemic, IMV was employed early, or at a lower supplemental oxygen delivery threshold, to manage ARDS, but its usage has been associated with several iatrogenic risks such as ventilator-associated pneumonia (Wicky et al., 2021) and barotrauma (Shrestha et al., 2022). Of interest to the present study is acute kidney injury (AKI), a critical condition that complicates ICU stays and is associated with mortality (Vemuri et al., 2022). As more data emerged and practitioners became more familiar with the pace of disease progression, guidelines shifted towards delaying intubation due to high mortality rates among mechanically ventilated patients and potential secondary complications. To navigate these treatment dynamics, we aim to elucidate the direct and indirect effects of IMV on patient mortality with respect to the AKI pathway. This application offers a compelling case for the use of our methods in understanding the effects of time-varying treatments and the causal pathways involved. We emphasize the phenomenon of “inconsistent mediation” (a situation in which the direct and indirect effects are qualitatively in opposite directions). In the current setting, this would involve (on average) IMV causing death through AKI while (on average) preventing death through other mechanisms. This dichotomy may be relevant for research and clinical practice, if interventions could be in place to mitigate IMV-induced AKI. While this is an active area of research, potential interventions include the administration of diuretics to modulate intravascular volume status (Grams et al., 2011; Glassford and Bellomo, 2011) and pharmacological options as explored by Pickkers et al. (2022). Finally, we also investigate the relationship between baseline covariates and the magnitude of these direct and indirect effects; such relationships can suggest biological pathways as well as potentially inform future treatments.
@@ 2.1 Notation
Our notation largely mirrors that of Díaz et al. (2021) and Díaz et al. (2023b). We consider data $X_{1},\ldots,X_{n}$, an i.i.d. sample from a distribution $\mathsf{P}$. The data are longitudinal, with $X=(L_{1},A_{1},Z_{1},M_{1},L_{2},\ldots,A_{\tau},Z_{\tau},\allowbreak M_{\tau},L_{\tau+1})$, where the variables $A$ are the exposure of interest, $L$ are covariates, $M$ are mediators, and $Z$ are intermediate confounders. An intermediate confounder is a variable affected by exposure and which affects the mediator and outcome; see VanderWeele et al. (2014). We let $Y=L_{\tau+1}$ denote the outcome of interest. In our illustrative application, $A$ will denote intubation status, $M$ is the occurence of AKI, $Y$ is mortality/survival, and $L$ and $Z$ are covariates or lab results, either at baseline or during the hospital stay. As in our illustrative application, we will assume that the mediator takes values in a finite set; the supports of all other variables are unrestricted.
@@ 2.2 Structural equation model and definition of causal effects
We consider causal effects defined within a nonparametric structural equation model, which consists of exogenous and endogenous variables and unknown deterministic functions governing the causal relations between them, together with independence assumptions on the exogenous variables (Pearl, 2000). We allow for arbitrary causal relations among observed variables that respect their time-ordering; this can be visualized in the saturated directed acyclic graph (DAG) given in Figure 1, reproduced from Díaz et al. (2023b).
@@
To formalize the dependence structure, we posit functions $f$ such that for each $t$, we have $A_{t}=f_{A,t}(H_{A,t},U_{A,t})$, $Z_{t}=f_{Z,t}(H_{Z,t},U_{Z,t})$, $M_{t}=f_{M,t}(H_{M,t},U_{M,t})$, and $L_{t}=f_{L,t}(H_{L,t},U_{L,t})$. The variables $U=(U_{A,t},U_{Z,t},U_{M,t},U_{L,t},U_{Y}:t\in\{1,\ldots,\tau\})$ are unobserved “exogenous” variables. In the general model, there are no restrictions on the distribution of $U$, but in Section 3, they will be subject to assumptions to enable identification of the estimands of interest.
@@
For interventions on the treatment variable, we will consider longitudinal modified treatment policies (LMTPs) (Díaz et al., 2021), which are hypothetical interventions defined in terms of a sequence of functions $\mathsf{d}=(\mathsf{d}_{1},\ldots,\mathsf{d}_{\tau})$, where $\mathsf{d}_{t}$ is a function $\mathsf{d}_{t}(a_{t},h_{A,t},\epsilon_{t})$ that depends on a treatment value $a_{t}$, a history value $h_{A,t}$, and possibly a randomizer $\epsilon_{t}$. (We will not consider such a randomizer in this paper.) The intervention is defined by removing the equation $A_{t}=f_{A_{t}}(H_{A,t},U_{A,t})$ sequentially (from $t=1$ to $t=\tau$) from the structural model, and replacing it with evaluations of the functions $\mathsf{d}_{t}$ as follows. At the first timepoint, the LMTP assigns the exposure as a new random variable $A_{1}^{\mathsf{d}}=\mathsf{d}_{t}(A_{1},H_{A,1},\epsilon_{1})$. This generates a counterfactual treatment $A_{2}(\bar{A}_{1}^{\mathsf{d}})=f_{A_{2}}(H_{A,2}(\bar{A}_{1}^{\mathsf{d}}),U_{A,1})$ that would have been observed if the intervention had been discontinued right before $t=2$. Then, the intervention at $t=2$ assigns treatment as $A_{2}^{\mathsf{d}}=\mathsf{d}_{2}(A_{2}(\bar{A}_{1}^{\mathsf{d}}),H_{A,2}(\bar{A}_{1}^{\mathsf{d}}),\epsilon_{2})$.
Across multiple timepoints, it is helpful to define the counterfactual history $H_{A,t}(\bar{A}_{t-1}^{\mathsf{d}})$ as the sequence of variables prior to $A_{t}$ that would have been observed under the LMTP given by $\bar{A}_{t-1}^{\mathsf{d}}$. In addition, $A_{t}(\bar{A}_{t-1}^{\mathsf{d}})=f_{A_{t}}(H_{A,t}(\bar{A}_{t-1}^{\mathsf{d}}),U_{A,t})$, termed the “natural value of treatment” (Richardson and Robins, 2013; Young et al., 2014), “represents the value of treatment that would have been observed at time $t$ under an intervention carried out up until time $t-1$ but discontinued thereafter” (Díaz et al., 2021). The intervention at time $t$ is given by $A_{t}^{\mathsf{d}}=\mathsf{d}_{t}(A_{t}(\bar{A}_{t-1}^{\mathsf{d}}),H_{A,t}(\bar{A}_{t-1}^{\mathsf{d}}),\epsilon_{t})$.
@@
The intervention on $\bar{M}$ is defined in terms of a stochastic intervention. Specifically, let $M_{t}(\bar{A}^{\mathsf{d}})=f_{M,t}(A_{t}^{\mathsf{d}},H_{M,t}(\bar{A}_{t}^{\mathsf{d}}),U_{M,t})$ denote the mediator at time $t$ observed under the above LMTP, and let $\bar{J}(\bar{A}^{\mathsf{d}})$ denote a random draw from the distribution of $\bar{M}(\bar{A}^{\mathsf{d}})$, potentially conditioning on baseline variables $L_{1}$. We consider interventions where $\bar{M}$ is replaced by this random draw.
@@
A causal “effect” is typically conceived as some difference in outcome distributions under different interventions. To simplify notation throughout, we will consider two user-specified, fixed interventions $\mathsf{d}^{\star}=(\mathsf{d}^{\star}_{1},\ldots,\mathsf{d}^{\star}_{\tau})$ and $\mathsf{d}^{\prime}=(\mathsf{d}^{\prime}_{1},\ldots,\mathsf{d}^{\prime}_{\tau})$, though the generalization of the theory to arbitrary $\mathsf{d}$ is immediate. We define the randomized modified treatment policy effect as $\mathsf{E}[Y(\bar{A}^{\mathsf{d}^{\prime}},\bar{J}(\bar{A}^{\mathsf{d}^{\prime}}))-Y(\bar{A}^{\mathsf{d}^{\star}},\bar{J}(\bar{A}^{\mathsf{d}^{\star}}))]$, where $Y(\bar{A},\bar{J})$ denotes the potential outcome of $Y$ where the treatment is intervened on to equal $\bar{A}$, and the mediator is intervened on to equal $\bar{J}$. We decompose this into an interventional direct effect
$DE=\mathsf{E}[Y(\bar{A}^{\mathsf{d}^{\prime}},\bar{J}(\bar{A}^{\mathsf{d}^{\star}}))-Y(\bar{A}^{\mathsf{d}^{\star}},\bar{J}(\bar{A}^{\mathsf{d}^{\star}}))]$ (1)
@@
$IE=\mathsf{E}[Y(\bar{A}^{\mathsf{d}^{\prime}},\bar{J}(\bar{A}^{\mathsf{d}^{\prime}}))-Y(\bar{A}^{\mathsf{d}^{\prime}},\bar{J}(\bar{A}^{\mathsf{d}^{\star}}))],$ (2)
which sum to the “overall” or “total” effect comparing $\mathsf{d}^{\prime}$ to $\mathsf{d}^{\star}$. Thus, the task at hand is to estimate expressions of the form $\mathsf{E}[Y(\bar{A}^{\mathsf{d}^{\prime}},\bar{J}(\bar{A}^{\mathsf{d}^{\star}}))]$, for general $\mathsf{d}^{\star}$ and $\mathsf{d}^{\prime}$.
For more details on the interpretation of the interventional effect, see Section S2 of the supplement, which also describes the nuance of the direct/indirect distinction for longitudinal data.
@@ 3.1 Assumptions and sequential regression
In what follows, it will be necessary to reference variable histories under interventions. For example, we let $H_{L,t}^{\prime}=(\bar{L}_{t-1},\bar{M}_{t-1},\bar{Z}_{t-1},\bar{A}_{t-1}^{\mathsf{d}^{\prime}})$. Intervened histories for other variables are defined analogously, as are counterfactual histories under the other intervention $\mathsf{d}^{\star}$, denoted with $H^{\star}$. Similarly, to ease notation, we let $A_{t}^{\prime}=\mathsf{d}_{t}^{\prime}(A_{t},H_{A,t},\epsilon_{t})$, with $A_{t}^{\star}$ defined similarly.
To identify the parameter, $\mathsf{E}[Y(\bar{A}^{\mathsf{d}^{\prime}},\bar{J}(\bar{A}^{\mathsf{d}^{\star}}))]$, the following assumptions, which combine elements of the identification assumptions from Díaz et al. (2021) and Díaz et al. (2023b), are sufficient. Note that 1 involves future histories rather than single timepoints only.
1Conditional exchangeability of treatment and mediator.
@@
    $U_{A,t}\mbox{$\perp\!\!\!\perp$}(\underline{U}_{Z,t},\underline{U}_{M,t},\underline{U}_{L,{t+1}},\underline{U}_{A,t+1})\mid H_{A,t}$ for all $t\in\{1,\ldots,\tau\}$;
@@
    $U_{M,t}\mbox{$\perp\!\!\!\perp$}(\underline{U}_{L,{t+1}},\underline{U}_{A,t+1},\underline{U}_{Z,{t+1}})\mid H_{M,t}$ for all $t\in\{1,\ldots,\tau\}$.
2Positivity of treatment and mediator assignment.
@@
    If $(a_{t},h_{A,t})\in\mathop{\mathrm{supp}}\{A_{t},H_{A,t}\}$ then $(\mathsf{d}^{\star}(a_{t},h_{A,t}),h_{A,t})\in\mathop{\mathrm{supp}}\{A_{t},H_{A,t}\}$ for $t\in\{1,\ldots,\tau\}$, and similarly for $\mathsf{d}^{\prime}$;
@@
3Piecewise-smooth invertibililty of treatment modification.
If $A_{t}$ is not a discrete random variable, the conditional support of $A_{t}$ given $H_{A,t}=h_{A,t}$ admits a partition into subintervals where the restrictions of $\mathsf{d}^{\prime}(\cdot,h_{a,t})$ and $\mathsf{d}^{\star}(\cdot,h_{a,t})$ to each subinterval have differentiable inverse functions.
Our identification strategy makes use of a representation of longitudinal causal estimands by means of “sequential regression” as proposed by Bang and Robins (2005) and further investigated by van der Laan and Gruber (2012); Luedtke et al. (2017), and Rotnitzky et al. (2017). This approach considers a recursive regression of outcomes (or pseudo-outcomes) onto previous variables, starting at $t=\tau$ and continuing backward to $t=0$. As in Díaz et al. (2023b), we split the recursion over the two non-interventional variables $L$ and $Z$ as follows. We consider fixed, pre-specified interventions $\mathsf{d}^{\prime}$, $\mathsf{d}^{\star}$, and $\bar{m}$. Starting with, $\mathsf{Q}_{Z,\tau+1}=Y$, define
@@
When the meaning is clear, noting that the only non-random component of the functions $\mathsf{Q}$ are the user-specified $\mathsf{d}^{\prime},\mathsf{d}^{\star}$, and $\bar{m}$, we may use the abbreviations $\mathsf{Q}_{L,t}^{\prime}(\underaccent{\bar}{m}_{t})=\mathsf{Q}_{L,t}(\bar{H}_{M,t},\underaccent{\bar}{m}_{t})$, $\mathsf{Q}_{Z,t}^{\prime}(\underaccent{\bar}{m}_{t})=\mathsf{Q}_{Z,t}(A^{{}^{\prime}}_{t},\bar{H}_{A,t},\underaccent{\bar}{m}_{t})$, and $\mathsf{Q}_{M,t}^{\star}(\underaccent{\bar}{m}_{t})=\mathsf{Q}_{M,t}(A_{t}^{\star},\bar{H}_{A,t},\underaccent{\bar}{m}_{t})$. We give the following identification result, which is a generalization of Theorem 1 of Díaz et al. (2023b).
@@
Under Assumptions 1 and 2, the interventional parameter $\theta=\mathsf{E}[Y(\bar{A}^{\mathsf{d}^{\prime}},\bar{J}(\bar{A}^{\mathsf{d}^{\star}}))]$ is identified as $\theta=\sum_{\bar{m}}\varphi(\bar{m})\lambda(\bar{m})$, where $\varphi(\bar{m})=\mathsf{Q}_{L,0}^{\prime}(\bar{m})$ and $\lambda(\bar{m})=\mathsf{Q}_{M,0}^{\star}(\bar{m})$.
Proof See Section S2 of the supplement. ∎